OpenStudy (anonymous):

The midpoint of is (5, –11). The coordinates of one endpoint are U(3, 5). Find the coordinates of endpoint V. But please don't just give me the answer. I need help!

OpenStudy (anonymous):

let V=(h,k) h+3=5*2; h=7; k+5=-11*2 k=-27 V=(7,-27)

OpenStudy (lalaly):

\[\large{\underline{midpoint(x, y)}}= (\frac{x_u+x_v}{2},\frac{y_u+y_v}{2})\]\[\large{5=\frac{3+x_v}{2}}\]\[\large{x_v=7}\] \[\large{-11=\frac{5+y_v}{2}}\]\[\large{y_v=-27}\] \[\huge{V=(7,-27)}\]

OpenStudy (anonymous):